OpenStudy (anonymous):

One factor for 2x2 + 16x – 18 is x + 9 ??

OpenStudy (anonymous):

Put x = -9 in the given expression and check whether it is coming zero or not?

OpenStudy (anonymous):

\[2x^2 + 16x -18\] Here put x = -9 and tell whether you get zero or not?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Then this is the factor of that expression.. See if I say that (x - 9) is the factor of expression x^2 - 81, then firstly put x - 9 = 0 so x = 9 Put this value of x in that expression and if it comes zero then (x - 9) is the factor of that expression... That is why I put x + 9 = 0 x = -9 and you substitute x = -9 to get zero..